Choose the ordered pair that is a solution to the system of equations.
3x - y = 9
2x + y = 6

Answer by mangopeeler07(462)

3x - y = 9
2x + y = 6

Add them
5x=15
x=3

Plug that in
2(3)+y=6
6+y=6
y=0

(3,0)